OpenStudy (themockingjay):

How long would it take to travel 37 km at a speed of 6000 m/hr?

OpenStudy (lord_box):

First you want to convert 6000 m/hr into km/hr, for this, we use dimensional analysis:\[\frac{ 6000m }{ 1hr } *\frac{ 1km }{ 1000km }\]

OpenStudy (themockingjay):

6000/1000? @Lord_Box

OpenStudy (lord_box):

Yes, and the units should cancel out to m/hr (the bottom km should me m, sorry)

OpenStudy (lord_box):

\[\frac{ 37km }{ 1 }*\frac{ 1hr }{ 6km }\]

OpenStudy (themockingjay):

how did you get that? @lord_box

OpenStudy (lord_box):

On the first step, we converted 6000m/hr to 6km/hr

OpenStudy (lord_box):

And so we can reason that if 6 km takes 1 hour, then 37 km should take 37/6 hours.

OpenStudy (themockingjay):

So the answer in 37/6?

OpenStudy (themockingjay):

@Lord_Box

OpenStudy (lord_box):

Yes, and make sure you include your unit hr.